# Rethinking Discrete Speech Representation Tokens for Accent Generation

### v1

@@ Abstract
Discrete Speech Representation Tokens (DSRTs) have become a foundational component in speech generation. While prior work has extensively studied phonetic and speaker information in DSRTs, how accent information is encoded in DSRTs remains largely unexplored. In this paper, we present the first systematic investigation of accent information in DSRTs. We propose a unified evaluation framework that measures both accessibility of accent information via a novel Accent ABX task and recoverability via cross-accent Voice Conversion (VC) resynthesis. Using this framework, we analyse DSRTs derived from a variety of speech encoders. Our results reveal that accent information is substantially reduced when ASR supervision is used to fine-tune the encoder, but cannot be effectively disentangled from phonetic and speaker information through naive codebook size reduction. Based on these findings, we propose new content-only and content-accent DSRTs that significantly outperform existing designs in controllable accent generation. Our work highlights the importance of accent-aware evaluation and provides practical guidance for designing DSRTs for accent-controlled speech generation.
Rethinking Discrete Speech Representation Tokens for Accent Generation
Jinzuomu Zhong1thanks: Equal contribution., Yi Wang [^1] footnotemark: 1, Korin Richmond1, Peter Bell1, 1Centre for Speech Technology Research, University of Edinburgh Correspondence: {jinzuomu.zhong, korin.richmond}@ed.ac.uk
@@ 1 Introduction
Inspired by the tokenisation schemes employed by text-based Large Language Models (LLMs), discrete speech tokenisation has emerged as an active research area for bridging speech and LLMs. By compressing speech signals or learned continuous representations into a finite set of discrete units, discrete speech tokens enable efficient storage of complex acoustic information while remaining inherently compatible with LLM architectures. As a result, they are rapidly becoming established as a foundational representation across a wide range of speech tasks (Guo et al., 2025).
@@
Discrete speech representation tokens (DSRTs), quantised from speech representations, have consequently seen huge success in various speech generation tasks, including Speech Language Models (SpeechLMs) (Cui et al., 2025), Zero-Shot Text-to-Speech (ZS-TTS) (Kharitonov et al., 2023), Speech-to-Speech Translation (S2ST) (Lee et al., 2022) and full-duplex dialogue systems (Nguyen et al., 2023).
@@
Despite the significant progress brought by DSRTs, the representation of a speaker’s accent is largely overlooked in the design, evaluation, and application of the tokens. Prior user perception studies have demonstrated a similarity–attraction effect, whereby listeners prefer accents similar to their own in spoken interaction (Dahlbäck et al., 2007). Unfortunately, ZS-TTS systems are shown to hallucinate accents that differ from those of the reference speaker (Zhong et al., 2025b), while spoken dialogue systems still lack broad adoption of benchmarks or evaluation settings that consider appropriate, accent-controlled speech generation (Cheng et al., 2025).
@@
Although some recent zero-shot TTS (ZS-TTS) systems have demonstrated some ability to mimic or control the accent of generated speech using discrete speech representation tokens (DSRTs), how much accent information is encoded in these tokens remains unexplored (Du et al., 2024; Zhang et al., 2025; Wang et al., 2025). Existing claims – such as that naive codebook size adjustment (Zhang et al., 2025) or ASR supervision (Du et al., 2024) can facilitate controlled accent generation – lack systematic investigation. To this end, we ask the following research questions: (1) How do different design choices in DSRTs influence the amount of accent information they encode? (2) How can these insights be leveraged to enable more controllable accent generation, e.g. in voice conversion (VC)?
@@
Existing DSRT evaluation frameworks focus on phonetic and speaker information and rarely consider accent (Polyak et al., 2021). We extend the investigation of accessibility to accent information, with an accent ABX method we proposed that evaluates the discriminability of representations for words uttered in different accents. Additionally, we propose to evaluate the recoverability of accent, speaker, and phonetic information by cross-accent VC, resynthesising from DSRTs from source speakers and speaker IDs from target speakers of different accents. Using the proposed framework, we observe the following key findings: (1) Accent information is most prominent in mid-early layers of HuBERT, different from speaker or phonetic information distribution. (2) Naive adjustment of codebook sizes provides only very limited disentanglement of accent, speaker and content information. (3) Predominant design of DSRTs for speech generation (quantising a later layer in a speech representation model, or using ASR supervison) discards most accent information.
Our main contributions are:
  • To the best of our knowledge, this is the first work to incorporate accent in evaluating DSRTs, and to apply ABX to a study of accent.
@@
  • Based on our findings, we propose quantisation schemes that are appropriate for accent-preserving VC – preserving source speaker accent – and accent-adaptive VC – adapting to target speaker accent – which achieve superior performance to existing approaches.
  • Evaluation pipeline and proposed DSRTs will be open-sourced in the future.
@@ 2.1 Discrete Speech Tokens
Discrete speech tokens can be largely categorised into two groups: semantic tokens [^1] A severe misnomer as these tokens contain primarily phonetic information and little semantic information (Wells et al., 2022; Choi et al., 2024). and acoustic tokens. We refer to semantic tokens as Discrete Speech Representation Tokens (DSRTs) here, naming them strictly by how these tokens are designed and trained, rather than what information some researchers assume them to possess. DSRTs are tokens obtained by quantising learned speech representations, such as HuBERT (Hsu et al., 2021) and w2v-BERT (Chung et al., 2021); while acoustic tokens are obtained by quantising waveforms. Common quantisation techniques include k-means, Vector Quantisation (VQ) (van den Oord et al., 2017), Residual Vector Quantisation (RVQ) (Zeghidour et al., 2021), and Finite-State Quantisation (FSQ) (Mentzer et al., 2024), with k-means, VQ, and FSQ primarily used for DSRTs, VQ and RVQ primarily used for acoustic tokens.
@@ 2.2 Investigation of Speech Representations
In terms of of continuous speech representations, most previous work has focused on accessibility, finding that phonetic information peaks in intermediate layers, while acoustic information dominates earlier layers (Choi et al., 2024; Pasad et al., 2021). After finetuning on the ASR task, it has been observed deeper layers become dominated by task-specific information, exhibiting higher phonetic discrimination.
In the investigation of DSRTs, Wells et al. (2022) find that DSRTs obtained by k-means on hubert-base-ls960 correspond to sub-phonetic events. Yeh and Tang (2024), using their information completeness and accessibility framework, found that DSRTs obtained from hubert-base-ls960 layer 4 contain more accessible pitch and speaker information, but less accessible phone information, than layer 9. Phonetic and speaker information in both layers consistently decrease with lower bitrates of RVQ codes, refuting claims that content and speaker disentanglement can be achieved by VQ.
@@ 2.3 DSRTs in ZS-TTS and SpeechLMs
In ZS-TTS, researchers found that DSRTs provide more accessible phonetic information than acoustic tokens do, while suffering information loss such as timbre and acoustic environment (Wang et al., 2025). Motivated by such findings, many state-of-the-art (SOTA) models utilise a hierarchical approach, first predicting DSRTs from text or source speech and then predicting acoustic tokens from DSRTs (Du et al., 2024; Lee et al., 2025; Wang et al., 2025; Zhang et al., 2025). In SpeechLMs, researchers seek to combine DSRTs with acoustic tokens for both accessible and complete/recoverable information, such as SpeechTokenizer (Zhang et al., 2024) and Mimi tokeniser (Défossez et al., 2024).
When designing such DSRTs, ZS-TTS researchers have claimed that are unverified or even contradicted by speech representation investigations. Zhang et al. (2025) choose layer 18 of hubert-large-ll60k and claim that reducing the codebook size in VQ leads to natural disentanglement of speaker, style (accent and emotion by their definition) and content, with a codebook size of 32 containing only content, and a codebook size of 8,192 containing only content and style. This claim not only contradicts the findings of Yeh and Tang (2024) that VQ or k-means with 1024 codebook size still retain significant speaker information, but also lacks verification in terms of accent and emotion. Du et al. (2024) propose to use “supervised semantic token”, obtained by injecting FSQ in an internal ASR model encoder; how ASR pretraining affects the information in discrete tokens remains unclear.
@@ 2.4 Accent Control and Generation
Unfortunately, ZS-TTS speech generation systems suffer accent hallucination, whereby they generate a hallucinated synthetic accent that deviates from the reference or prompt speech (Zhong et al., 2025b). Recent systems that have been specially built for accent generation have not yet utilised DSRTs along with powerful LLM architectures (Zhong et al., 2025b; Xinyuan et al., 2025). General ZS-TTS systems are reported to have certain accent control capabilities, but there has been little or no investigation about accent information in DSRTs. Understanding how accent information is encoded could help build more inclusive ZS-TTS systems for diverse accents.
@@ 2.5 ABX Evaluation
[…]
Previous work introduced Minimal Pair ABX (MP-ABX) (Schatz et al., 2013), where instances differ only in a central phonetic unit while sharing the same phonetic context. Prior studies (Borsos et al., 2023; Wang et al., 2025; Zhang et al., 2025) have used this metric to assess the phonetic discriminability of DSRT tokens. In addition, ABX has been extended to probe speaker-related information by defining triplets that contrast speaker identity under controlled linguistic content (de Seyssel et al., 2022). Nevertheless, to the best of our knowledge, no prior work has systematically employed machine ABX to reveal accent discriminability in speech representations or DSRTs.
@@ 3 Method
Inspired by the perspective of separately measuring information completeness and accessibility Yeh and Tang (2024), we focus on evaluating DSRTs from both aspects, but in the context of accent generation. Since completeness cannot be measured by resynthesis, we introduce recoverability as a task-grounded measure of how much accent, speaker, and phonetic information encoded in DSRTs can be recovered in resynthesised speech. Together with accessibility, we propose a pipeline that evaluates DSRTs from both synthesis-facing and representation-facing perspectives.
Figure 1 provides an overview of the proposed pipeline. First, we obtain DSRTs from several commonly used speech representations using RepCodec Huang et al. (2024) with Vector Quantisation (VQ) van den Oord et al. (2017) as the quantiser (see Section 3.1). For each DSRT configuration, we then train a unit-to-speech resynthesis model using HiFiGAN Polyak et al. (2021). To assess information recoverability, we infer unit-to-speech models with DSRTs from a source speaker and speaker ID from a target speaker with a different accent, thereby conducting cross-accent VC. The generated speech is evaluated using a combination of objective metrics and subjective listening tests (see Section 3.2). Finally, to assess information accessibility, we directly probe the DSRTs using a range of ABX setups (see Section 3.3).
@@ 3.1 Discrete Speech Representation Tokens
We choose three speech representation models to obtain Discrete Speech Representation Tokens (DSRTs): HuBERT; HuBERT finetuned for ASR (HuBERT-ft); and Whisper. We include HuBERT Hsu et al. (2021) as it provides the most widely used speech representations in various speech tasks such as SpeechLM and TTS Guo et al. (2025); Cui et al. (2025). Recently, there has been a growing trend toward using ASR-based speech representations for obtaining DSRTs. Therefore, we include HuBERT-ft and Whisper as two representative ASR-based representations, with Encoder-only and Encoder-Decoder architectures, respectively.
Following RepCodec Huang et al. (2024), we use VQ-VAE van den Oord et al. (2017) to discretise the speech representations. The model consists of three modules: Encoder and Decoder, which are both convolution layers with residual paths, and a Vector Quantisation module that quantises latent representations from Encoder output into a series of discrete tokens. The model is trained to reconstruct the speech representation, with additional Exponential Moving Average (EMA) optimisation to gradually update the codebook. For model details, we refer readers to the original paper Huang et al. (2024).
@@ 3.2 Cross-Accent Voice Conversion
After extracting DSRTs, we train a unit-to-speech HiFiGAN model for each DSRT configuration, following Polyak et al. (2021). In prior work, information recoverability is typically evaluated using resynthesis or Voice Conversion (VC) on General American English datasets, with evaluation focusing primarily on intelligibility and speaker similarity, while accent is largely ignored.
We train unit-to-speech models on data covering multiple accents to assess the generalisability of DSRTs across accents. We then perform cross-accent VC in inference, by conditioning the model on DSRTs from a source speaker and a target speaker ID with a different accent. Afterwards, during evaluation, following Zhong et al. (2025a), we explicitly assess accent, speaker, and phonetic similarity in the converted speech. Utterance-level embeddings from Accent Identification (AID) and Speaker Verification (SV) models are used to calculate cosine similarity, which serves as a proxy for perceived accent and speaker similarity. Phonetic Posteriorgrams (PPGs) are extracted, aligned, and then used to calculate phonetic distances as a proxy for perceived phonetic similarity, and to some degree, intelligibility.

### v2

# Zhong Wang Richmond Bell

@@ Abstract
Discrete Speech Representation Tokens (DSRTs) have become a foundational component in speech generation. While prior work has extensively studied phonetic and speaker information in DSRTs, how accent information is encoded in DSRTs remains largely unexplored. In this paper, we present the first systematic investigation of accent information in DSRTs. We propose a unified evaluation framework that measures both accessibility of accent information via a novel Accent ABX task and recoverability via cross-accent Voice Conversion (VC) resynthesis. Using this framework, we analyse DSRTs derived from several widely used speech representations. Our results reveal that: (1) choice of layers has the most significant impact on retaining accent information, (2) accent information is substantially reduced by ASR supervision; (3) naive codebook size reduction cannot effectively disentangle accent from phonetic and speaker information.
keywords:
accent, discrete speech representation, voice conversion, accent conversion
@@ 1 Introduction
Inspired by the tokenisation schemes employed by text-based Large Language Models (LLMs), discrete speech tokenisation has emerged as an active research area for bridging speech and LLMs. By compressing speech signals or learned continuous representations into a finite set of discrete units, discrete speech tokens enable efficient storage of complex acoustic information while remaining inherently compatible with LLM architectures. As a result, they are rapidly becoming established as a foundational representation across a wide range of speech tasks [guo2025recent].
@@
Directly compressing raw speech signals, however, tends to produce a representation that preserves abundant low-level acoustic detail, making phonetic content difficult to access. Such an approach works well for waveform reconstruction, but places a heavier modelling burden on downstream speech generation tasks [wang2025maskgct]. To address this, discrete speech representation tokens (DSRTs), which are quantised from speech representations, are often used. DSRTs have seen huge success in various speech generation tasks, including Speech Language Models (SpeechLMs) [cui2025recent], Zero-Shot Text-to-Speech (ZS-TTS) [kharitonov2023speak], Speech-to-Speech Translation (S2ST) [lee2022direct] and full-duplex dialogue systems [nguyen2023dgslm].
@@
Despite the significant progress brought by DSRTs, the representation of a speaker's accent is largely overlooked in the design, evaluation, and application of the tokens. Prior user perception studies have demonstrated a similarity–attraction effect, whereby listeners prefer accents similar to their own in spoken interaction [dahlback2007similarity]. Unfortunately, ZS-TTS systems are shown to hallucinate accents that differ from those of the reference speakers [zhong2025accentbox], while spoken dialogue systems still lack broad adoption of benchmarks or evaluation settings that consider appropriate, accent-controlled speech generation [cheng2025voxdialogue].
@@
Although some recent ZS-TTS systems have demonstrated some ability to mimic or control the accent of generated speech using DSRTs, how much accent information is encoded in these tokens remains unexplored [du2024cosyvoice, zhang2025vevo, wang2025maskgct]. Existing claims – such as that naive codebook size adjustment [zhang2025vevo] or ASR supervision [du2024cosyvoice] can facilitate accent control and generation – lack systematic investigation. Crucially, the accent information actually preserved in these DSRTs remains unquantified, leaving it unclear whether the observed accent generation capabilities in ZS-TTS stem from the representations themselves or are merely a byproduct of large-scale pretraining. To this end, we ask the following research questions: (1) How do different design choices in DSRTs influence the amount of accent information they encode? (2) How can these insights be leveraged to enable more controllable accent generation, e.g. in voice conversion (VC)?
@@
To answer the research questions, we propose a framework for evaluating DSRTs from both the recoverability and accessibility perspectives, in terms of accent, speaker, and phonetic information encoded. Existing DSRT evaluation frameworks focus on phonetic and speaker information and rarely consider accent [polyak2021speech]. To the best of our knowledge, this is the first work to incorporate accent in evaluating DSRTs, and to apply ABX to the study of accent. Specifically, we extend the existing investigation of accessibility to accent information, with a novel accent ABX method that evaluates the discriminability of representations for words uttered in different accents. Additionally, we propose to evaluate the recoverability of accent, speaker, and phonetic information by cross-accent VC, resynthesising with DSRTs from source speakers and speaker IDs from target speakers of different accents as input.
Using this proposed framework, we observe the following key findings: (1) Accent information is most prominent in mid-early layers of HuBERT, different from speaker or phonetic information distribution. (2) Naive adjustment of codebook sizes provides only very limited disentanglement of accent, speaker and content information. (3) Predominant design of DSRTs for speech generation (quantising a later layer in a speech representation model, or using ASR supervision) discards most accent information. These findings challenge existing claims using naive codebook size adjustment [zhang2025vevo] or ASR supervision [du2024cosyvoice] for accent control and generation.
@@
Finally, based on our findings, we propose DSRT design choices that are appropriate for both accent-preserving VC – preserving source speaker accent, and accent-adaptive VC – adapting to target speaker accent. Both objective and subjective evaluation show superior performance to existing approaches. We will release links to code and demo in the final version.
@@ 2.1 Discrete Speech Tokens
Discrete speech tokens can be largely categorised into two groups: semantic tokens [^1] A severe misnomer as these tokens contain primarily phonetic information and little semantic information [wells2022phonetic, choi2024self]. and acoustic tokens. We refer to semantic tokens as Discrete Speech Representation Tokens (DSRTs) here, naming them strictly by how these tokens are designed and trained, rather than what information some researchers assume them to possess. DSRTs are tokens obtained by quantising learned speech representations, such as HuBERT [hsu2021hubert] and w2v-BERT [chung2021w2vbert]; while acoustic tokens are obtained by directly quantising waveforms. Common quantisation techniques include k-means, Vector Quantisation (VQ) [vandenoord2017neural], Residual Vector Quantisation (RVQ) [zeghidour2021soundstream], and Finite-State Quantisation (FSQ) [mentzer2024finite], with k-means, VQ, and FSQ primarily used for DSRTs, VQ and RVQ primarily used for acoustic tokens.
@@ 2.2 Investigation of Speech Representations
In terms of of continuous speech representations, most previous work has focused on accessibility, finding that phonetic information peaks in intermediate layers, while acoustic information dominates earlier layers [choi2024self, pasad2021layer]. After finetuning on the ASR task, it has been observed deeper layers become dominated by task-specific information, exhibiting higher phonetic discrimination.
In the investigation of DSRTs, Wells et al. [wells2022phonetic] find that DSRTs obtained by k-means on hubert-base-ls960 correspond to sub-phonetic events. Yeh et al. [yeh2024estimating], using their information completeness and accessibility framework, found that DSRTs obtained from hubert-base-ls960 layer 4 contain more accessible pitch and speaker information, but less accessible phone information, than layer 9. Phonetic and speaker information in both layers consistently decrease with lower bitrates of RVQ codes, refuting claims that content and speaker disentanglement can be achieved by VQ.
@@ 2.3 DSRTs in ZS-TTS and SpeechLMs
In ZS-TTS, researchers found that DSRTs provide more accessible phonetic information than acoustic tokens do, while suffering information loss such as timbre and acoustic environment [wang2025maskgct]. Motivated by such findings, many state-of-the-art (SOTA) models utilise a hierarchical approach, first predicting DSRTs from text or source speech and then predicting acoustic tokens from DSRTs [du2024cosyvoice, lee2025hierspeechpp, wang2025maskgct, zhang2025vevo]. In SpeechLMs, researchers seek to combine DSRTs with acoustic tokens for both accessible and complete/recoverable information, such as SpeechTokenizer [zhang2024speechtokenizer] and Mimi tokeniser [defossez2024moshi].
When designing such DSRTs, ZS-TTS researchers have introduced some claims that are either unverified or contradicted by speech representation investigations. Vevo [zhang2025vevo] chooses layer 18 of hubert-large-ll60k and claims that reducing the codebook size in VQ leads to natural disentanglement of speaker, style (accent and emotion by their definition) and content, with a codebook size of 32 containing only content, and a codebook size of 8,192 containing only content and style. This claim not only contradicts the findings in [yeh2024estimating] that VQ or k-means with 1024 codebook size still retain significant speaker information, but also lacks verification in terms of accent and emotion. CosyVoice [du2024cosyvoice] proposes to use ``supervised semantic token'', obtained by injecting FSQ in an internal ASR model encoder, for content consistency in the generated speech; how ASR pretraining affects accent information in discrete tokens remains unclear.
@@ 2.4 Accent Control and Generation
Unfortunately, ZS-TTS speech generation systems suffer accent hallucination, whereby they generate a hallucinated synthetic accent that deviates from the reference or prompt speech [zhong2025accentbox]. Recent systems that have been specially built for accent generation have not yet utilised DSRTs along with powerful LLM architectures [zhong2025accentbox, xinyuan2025scalable]. General ZS-TTS systems are reported to have certain accent control capabilities, but there has been little or no investigation about accent information in DSRTs. Understanding how accent information is encoded could help build more inclusive ZS-TTS systems for diverse accents.
@@ 2.5 ABX Evaluation
The ABX error rate is a distance-based, model-agnostic evaluation metric, assessing whether the representations pull data of the same category (e.g. sample $x=$ [bIt], another sample $a=$ [bIt], of the same triphone category) closer and push dissimilar data (e.g. $x=$ [bIt], $b=$ [bi:t]) farther in representational space.
Previous work introduced Minimal Pair ABX (MP-ABX) [schatz2013evaluating], where instances differ only in a central phonetic unit while sharing the same phonetic context. Prior studies [wang2025maskgct, borsos2023audiolm, zhang2025vevo] have used this metric to assess the phonetic discriminability of DSRT tokens. In addition, ABX has been extended to probe speaker-related information by defining triplets that contrast speaker identity under controlled linguistic content [de2022language]. Nevertheless, to the best of our knowledge, no prior work has systematically employed machine ABX to reveal accent discriminability in speech representations or DSRTs.
@@ 3 Method
Inspired by the perspective of separately measuring information completeness and accessibility [yeh2024estimating], we focus on evaluating DSRTs from both aspects, but in the context of accent generation. Since completeness cannot be measured by resynthesis, we introduce recoverability as a task-grounded measure of how much accent, speaker, and phonetic information encoded in DSRTs can be recovered in resynthesised speech. Together with accessibility, we propose a pipeline that evaluates DSRTs from both synthesis-facing and representation-facing perspectives.
Figure 1 provides an overview of the proposed pipeline. First, we obtain DSRTs from several commonly used speech representations using RepCodec [huang2024repcodec] with Vector Quantisation (VQ) [vandenoord2017neural] as the quantiser (see Section 3.1). For each DSRT configuration, we then train a unit-to-speech resynthesis model using HiFiGAN [polyak2021speech]. To assess information recoverability, we infer unit-to-speech models with DSRTs from a source speaker and speaker ID from a target speaker with a different accent, thereby conducting cross-accent VC. The generated speech is evaluated using a combination of objective metrics and subjective listening tests (see Section 3.2). Finally, to assess information accessibility, we directly probe the DSRTs using a range of ABX setups (see Section 3.3).
@@ 3.1 Discrete Speech Representation Tokens
We choose three speech representation models to obtain Discrete Speech Representation Tokens (DSRTs): HuBERT; HuBERT finetuned for ASR (HuBERT-ft); and Whisper. We include HuBERT [hsu2021hubert] as it provides the most widely used speech representations in various speech tasks such as SpeechLM and TTS [guo2025recent, cui2025recent]. Recently, there has been a growing trend toward using ASR-based speech representations for obtaining DSRTs. Therefore, we include HuBERT-ft and Whisper [radford2023robust] as two representative ASR-based representations, with Encoder-only and Encoder-Decoder architectures, respectively. All three speech representations investigated, despite being at least three years old, are still widely used in current speech LLMs, evidenced by their usage in speech tokenisers of 19 systems [cui2025recent, Table 1].
Following RepCodec [huang2024repcodec], we use VQ-VAE [vandenoord2017neural] to discretise the speech representations. The model consists of three modules: Encoder and Decoder, which are both convolution layers with residual paths, and a Vector Quantisation module that quantises latent representations from Encoder output into a series of discrete tokens. The model is trained to reconstruct the speech representation, with additional Exponential Moving Average (EMA) optimisation to gradually update the codebook. For model details, we refer readers to the original paper [huang2024repcodec].
@@ 3.2 Cross-Accent Voice Conversion
After extracting DSRTs, we train a unit-to-speech HiFiGAN model for each DSRT configuration, following [polyak2021speech]. In prior work, information recoverability is typically evaluated using resynthesis or Voice Conversion (VC) on General American English datasets, with evaluation focusing primarily on intelligibility and speaker similarity, while accent is largely ignored.
We train unit-to-speech models on data covering multiple accents to assess the generalisability of DSRTs across accents. We then perform cross-accent VC in inference, by conditioning the model on DSRTs from a source speaker and a target speaker ID with a different accent. Afterwards, during evaluation, following [zhong2025pairwise], we explicitly assess accent, speaker, and phonetic similarity in the converted speech, against ground truth reference clips. Utterance-level embeddings from Accent Identification (AID) and Speaker Verification (SV) models are used to calculate cosine similarity, which serves as a proxy for perceived accent and speaker similarity. Phonetic Posteriorgrams (PPGs) are extracted, aligned, and then used to calculate phonetic distances as a proxy for perceived phonetic similarity, and to some degree, intelligibility.